Each of the four independent situations below describes a sales-type lease in which annual lease payments of $14,000 are payable at the beginning of each year. Each is a finance lease for the lessee. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Situation
1 2 3 4
Lease term (years) 3 3 3 3
Asset’s useful life (years) 3 4 4 6
Lessor’s implicit rate (known by lessee) 12 % 12 % 12 % 12 %
Residual value:
Guaranteed by lessee 0 $ 5,600 $ 2,800 0
Unguaranteed 0 0 $ 2,800 $ 5,600
Purchase option:
After (years) none 2 3 3
Exercise price n/a $ 7,800 $ 1,800 $ 3,800
Reasonably certain? n/a no no yes

  
Determine the following amounts at the beginning of the lease: (Round your final answers to nearest whole dollar.)

Situation
1 2 3 4
A. The lessor’s:
1. Lease payments
2. Gross investment in the lease
3. Net investment in the lease
B. The lessee’s:
4. Lease payments
5. Right-of-use asset
6. Lease payable

PV charts can be accessed through google.
Answer #1

Situation

1

2

3

4

A.

The lessor’s

1. Lease payments

42000

47600

47600

45800

2. Gross investment in the lease

42000

47600

47600

45800

3. Net investment in the lease

41084

45070

45070

43789

B.

The lessee’s:

4. Lease payments

42000

42000

42000

45800

5. Right-of-use asset

41084

41084

41084

43789

6. Lease payable

41084

41084

41084

43789

14000*3 = 42000

(14000*3)+5600 = 47600

(14000*3)+2800+2800 = 47600

(14000*3)+3800 = 45800

PV value of annuity due @12% for 3 years = 2.93460

PV value of @ 12% for 3 years = 0.71178

14000*2.93460 = 41084

(14000*2.93460)+(5600*0.71178) = 45070

(14000*2.93460)+(3800*0.71178) = 43789